Grade 12Algebra

Multiply and express as a mixed fraction:
7 × 2 ¼ …....

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

First convert the given mixed fraction into improper fraction.
= 2 whole 1⁄4
= 9/4
Now,
= 7 × (9/4)
= 63/4
= 15 whole 3⁄4

Thanks